OpenStudy (anonymous):

Segment Addition: Can someone help me solve this problem? In a rowing race, the distance between the teams in first and second place is 5.7 meters. The distance between the teams in second and third place is one-third that distance. How much farther ahead is the team in first place than the team in third?

OpenStudy (bahrom7893):

Distance b/n 1st and 2nd: 5.7 Distance b/n 2nd and 3rd is 1/3 that, so : 5.7 * (1/3) = 1.9

OpenStudy (bahrom7893):

Total distance: 5.7 + 1.9 = 13.3 <- Dis b/n 1 and 3

OpenStudy (bahrom7893):

fan me if i helped, please

OpenStudy (anonymous):

okay thanks! can you please explain to me how you got this answer? : )

OpenStudy (bahrom7893):

Well you have given: Distance b/n 1st and 2nd is 5.7 and distance between 2nd and 3rd is 1/3 of 5.7, which means divide 5.7 by 3

OpenStudy (bahrom7893):

so distance b/n 2nd and 3rd is 5.7/3 = 1.9

OpenStudy (anonymous):

so you multiply then divide?

OpenStudy (anonymous):

or just divide?

OpenStudy (bahrom7893):

So from 1 to 2 there is 5.7 and from 2 to 3 there is 1.9. Which means from 1 to 3 there is 5.7+1.9 = 7.6

OpenStudy (bahrom7893):

no u just divide, and sorry my previous answer was a typo, its not 13.3, the answer is 7.6

OpenStudy (anonymous):

hmm im totally confused : /

OpenStudy (anonymous):

is this the first step --> 5.7 * (1/3) = 1.9 then you would add 5.7+1.9= 7.6?

OpenStudy (anonymous):

& you get 7.6 as your answer?

OpenStudy (bahrom7893):

yes

OpenStudy (anonymous):

thank youuu! : )

OpenStudy (bahrom7893):

np